OpenStudy (anonymous):

The value of y varies directly with the value of x, and y = –21 when x = 3. Which equations represent the relationship between y and x? Choose exactly two answers that are correct. A. B. C. D. PLEASE HELP!!!

OpenStudy (anonymous):

A. -7X=Y B. Y/X=-7 C. X/Y=7 D. XY=-7

OpenStudy (anonymous):

@matthew1286 @mary212129 @Majesty69

OpenStudy (anonymous):

A

OpenStudy (welshfella):

direst variation is y = kx where k is a constant here, the value of k = y/x = -21/3 = -7